# How Does Sequence Modeling Architecture Influence Base Capabilities of Pre-trained Language Models? Exploring Key Architecture Design Principles

### v1

@@ 1 Introduction
However, due to its quadratic time complexity, the self-attention mechanism has long been plagued by high computational costs in long-sequence scenarios. To improve the efficiency of sequence modeling architectures, numerous novel stateful sequence modeling architectures have recently emerged, such as Mamba [16, 11], RWKV [31, 32, 33], Gated DeltaNet [51], etc. These architectures primarily inherit stateful modeling mechanisms of linear attention [23] and linear RNNs [28], offering advantages in time and space efficiency over self-attention. Additionally, they introduce new mechanisms like data-dependent decay and delta rule to enhance the expressive power of linearized modeling.
@@
To address these questions, we first point out that existing research on sequence modeling architectures typically adopts the same mixed domain pre-training settings used in large model development. While beneficial for practical applications, these settings are detrimental to architecture analysis, as they turn base capability tests (e.g., language modeling) into in-distribution evaluations, failing to reveal differences in base capabilities during early pre-training stages. To address this, we propose a limited domain pre-training approach with out-of-distribution testing, successfully uncovering significant differences in base capabilities among architectures at an early stage.
Next, under this setting, we analyze the base capabilities of stateful sequence modeling architectures like Mamba, RWKV, Gated DeltaNet, etc. Our findings reveal that these architectures exhibit notable degradation in base capabilities compared to Transformer, confirming that stateful sequence modeling architectures suffer from deficiencies not only in specialized capabilities but also in base capabilities.
We then investigate the architecture factors that truly impact base capabilities. Through ablation studies on the Mamba family of architectures and analyses of common sequence modeling factors, we identify that mechanisms like data-dependent decay, convolution and position encoding only affect convergence speed rather than base capabilities. Conversely, we determine that full-sequence visibility, real relation calculation and non-uniform distribution are critical architecture factors influencing base capabilities. Based on this, we summarize the principle that "a sequence modeling architecture need possess full-sequence arbitrary selection capability" as the key design rule to avoid degradation in base capabilities.
@@
Through our analysis and experiments, we prove the effectiveness of the proposed sequence modeling architecture design principle, providing valuable insights for future architecture improvements and novel designs.
@@ 3.2.1 Full-Sequence Visibility
The first factor we identified as critical to base capabilities is window size. This observation was prompted by the significant difference in base capabilities between Transformer++ and Transformer++ (Window=256), suggesting that window size may be a key factor.
To investigate, we tested various window sizes, as shown in Figure 3(d). The results demonstrate that as the window size increases, both pre-training convergence speed and base capabilities improve significantly. Conversely, reducing the sequence context leads to gradual degradation in base capabilities. Thus, we conclude that Full-Sequence Visibility is an essential requirement for sequence modeling architectures.
@@ 5.3 Long Sequence and Architecture Combination
Time efficiency and long-sequence retrieval results are presented in Table 5.3. Top-1 Chunk (Approx.) achieved substantial speed improvements over Transformer++, reaching time efficiency levels comparable to Mamba-2. Top-1 Chunk (Exact) also showed speed gains, achieving time efficiency similar to RWKV-7. In long-sequence retrieval, results on S-NIAH [20] shown Top-1 Chunk Selection delivered performance close to Transformer++, while other stateful models suffered severe degradation.
@@ Appendix A Model Specifications and Pre-training Procedures
Transformer adopts the GPT [35] architecture, is based on FlashAttention-2111https://github.com/Dao-AILab/flash-attention [10], and has a hidden dimension of 768, 12 layers and 12 attention heads for the small-scale version; while the large-scale version has a hidden dimension of 2048, 24 layers and 16 attention heads. Transformer++ introduces Rotary Embedding [43], GeGLU [41] and RMSNorm [56] over the Transformer, also built using FlashAttention-2, with identical model specifications as the Transformer. Transformer++ (W=256) incorporates a window size of 256, retains the same model specification, and is also based on FlashAttention-2.
Based [2] is implemented according to the official open-source code222https://github.com/HazyResearch/based, with a hidden dimension of 768 and 15 layers for the small-scale version; among them, there are 3 layers of window attention with 12 attention heads and a window size of 128, and 3 layers of TaylorExp linear attention with 12 attention heads. Mamba-1 [16] is implemented based on the official open-source code333https://github.com/state-spaces/mamba, with a hidden dimension of 768, 24 layers and a state size of 16 for the small-scale version; the large-scale version has a hidden dimension of 2048, 48 layers and the same state size of 16. Mamba-2 [11] is implemented based on the official open-source code (URL is the same as Mamba-1), with a hidden dimension of 768, 24 layers and a state size of 128 for the small-scale version; the large-scale version has a hidden dimension of 2048, 48 layers and the same state size of 128. RWKV-5/6/7 [32, 33] are implemented based on the official open-source code444https://github.com/BlinkDL/RWKV-LM, with a hidden dimension of 768, 12 layers and 12 attention heads for the small-scale versions. DeltaNet [40] is based on Flash Linear Attention555https://github.com/fla-org/flash-linear-attention [52], with a hidden dimension of 768, 12 layers and 8 attention heads for the small-scale version. Gated DeltaNet [51] is also based on Flash Linear Attention, with a hidden dimension of 768, 12 layers, 6 attention heads and a head dimension of 96 for the small-scale version.
Top-1 Element is implemented using PyTorch666https://pytorch.org/ and the transformers777https://github.com/huggingface/transformers library, matching the model specifications of the Transformer++. Top-1 Chunk (Exact/Approx.) shares the same model specifications as Transformer++, with a chunk size of 128. Its sequence modeling part is implemented using Triton888https://github.com/triton-lang/triton [46], and its offset and index CUDA kernel is modified from the fastmoe999https://github.com/laekov/fastmoe [18].
In addition, all of the above models use the Mixtral [22] vocabulary, which contains 32,000 tokens. Large-scale models involved in the training process were trained using the DeepSpeed101010https://github.com/deepspeedai/DeepSpeed open-source project. All models involving Rotary Embedding have a base of 10,000.
@@
Transformer++ is based on FlashAttention-2, while Transformer++ (W=256) uses FlexAttention111111https://pytorch.org/blog/flexattention/ [14]. The implementations of Top-1 Chunk (Exact) and Top-1 Chunk (Approx.) follow those described in the previous configuration. All four models have a hidden dimension of 768, 12 layers and 12 attention heads, with a chunk size of 128 for the two Top-1 Chunk models. Additionally, RWKV-7 was implemented using Flash Linear Attention when the sequence length was set to 100k to improve time efficiency. Gradient checkpointing was applied to reduce GPU memory consumption across all models. All models use the Mixtral vocabulary with a size of 32,000 tokens. Models incorporating Rotary Embedding maintain a base of 1,000,000.
@@ Appendix B Evaluation Procedures
For few-shot learning evaluation, general tasks include MMLU [19], OpenBookQA [29], ARC Easy & Challenge [9], BoolQ [8], RACE [24], SIQA [39], SCIQ [48], HellaSwag [55], COPA [37], PIQA [5], WinoGrande [38] and Winograd [25]; retrieval tasks include SWDE [26], SQuAD [36] and FDA [3]. Among these, MMLU uses a 5-shot setting, while all others are evaluated in a 0-shot setting. To ensure the stability and reproducibility of the evaluation results, we use the lm-evaluation-harness121212https://github.com/EleutherAI/lm-evaluation-harness [15] open-source evaluation framework for evaluation.
For time efficiency evaluation (Table 5.3), we compare the pre-training speed of different models with an input length of 100k tokens and a total batch size of 8. For inference speed comparison, we use an input length of 50k tokens and generate 5k tokens, with a total batch size of 64.

### v2

@@ 1 Introduction
However, due to its quadratic time complexity, the self-attention mechanism has long been plagued by high computational costs in long-sequence scenarios. To improve the efficiency of sequence modeling architectures, numerous novel stateful sequence modeling architectures have recently emerged, such as Mamba [16, 11], RWKV [31, 32, 33], Gated DeltaNet [51]. These architectures primarily inherit stateful modeling mechanisms of linear attention [23] and linear RNNs [28], offering advantages in time and space efficiency over self-attention. Additionally, they introduce new mechanisms like data-dependent decay and delta rule to enhance the expressive power of linearized modeling.
@@
To address these questions, we first point out that existing research on sequence modeling architectures typically adopts the same mixed domain pre-training settings used in large model development. While beneficial for practical applications, these settings are detrimental to architecture analysis, as they turn base capability tests (e.g., language modeling) into in-distribution evaluations, failing to reveal differences in base capabilities during early pre-training stages. To address this, we propose a limited domain pre-training approach, employing out-of-distribution language modeling performance to measure base capabilities, successfully uncovering significant differences in base capabilities among architectures at an early stage.
Next, under this setting, we analyze the base capabilities of stateful sequence modeling architectures like Mamba, RWKV, Gated DeltaNet. Our findings reveal that these architectures exhibit notable degradation in base capabilities compared to Transformer, confirming that stateful sequence modeling architectures suffer from deficiencies not only in specialized capabilities but also in base capabilities.
We then investigate the architecture factors that truly impact base capabilities. Through ablation studies on the Mamba family of architectures and analyses of common sequence modeling factors, we identify that mechanisms like data-dependent decay, convolution and position encoding only affect convergence speed rather than base capabilities. Conversely, we determine that full-sequence visibility, real relation calculation and non-uniform distribution are critical architecture factors influencing base capabilities. Based on this, we summarize the principle that "a sequence modeling architecture need to possess full-sequence arbitrary selection capability" as the key design rule to avoid degradation in base capabilities.
@@
Through our analysis and experiments, we prove the effectiveness of the proposed architecture design principle, providing valuable insights for future architecture designs.
@@ 3.2.1 Full-Sequence Visibility
The first factor we identified as critical to base capabilities is sliding window size. This observation was prompted by the significant difference in base capabilities between Transformer++ and Transformer++ (Window=256), suggesting that sliding window size may be a key factor.
To investigate, we tested various sliding window sizes, as shown in Figure 3(d). The results demonstrate that as the sliding window size increases, both pre-training convergence speed and base capabilities improve significantly. Conversely, reducing the sequence context leads to gradual degradation in base capabilities. Thus, we conclude that Full-Sequence Visibility is an essential requirement for sequence modeling architectures.
@@ 5.1.3 GPU Kernels Design
To achieve better time efficiency, we designed and implemented GPU kernels for the Top-1 Chunk Selection architecture. An overview is given in this section, with details in Appendix D .
@@ 5.3 Long Sequence and Architecture Combination
Time efficiency and long-sequence retrieval results are presented in Table 3. Top-1 Chunk (Approx.) achieved substantial speed improvements over Transformer++, reaching time efficiency levels comparable to Mamba-2. Top-1 Chunk (Exact) also showed speed gains, achieving time efficiency similar to RWKV-7. In long-sequence retrieval, results on S-NIAH [20] shown Top-1 Chunk Selection delivered performance close to Transformer++, while other stateful models suffered severe degradation.
@@ 7 Conclusion
Acknowledgments
This work was supported by the New Generation Artificial Intelligence-National Science and Technology Major Project 2023ZD0121100, the National Natural Science Foundation of China (NSFC) via grant 62441614 and 62176078.
@@ Appendix A Model Specifications and Pre-training Procedures
Transformer adopts the GPT [35] architecture, is based on FlashAttention-2222https://github.com/Dao-AILab/flash-attention [10], and has a hidden dimension of 768, 12 layers and 12 attention heads for the small-scale version; while the large-scale version has a hidden dimension of 2048, 24 layers and 16 attention heads. Transformer++ introduces Rotary Embedding [43], GeGLU [41] and RMSNorm [56] over the Transformer, also built using FlashAttention-2, with identical model specifications as the Transformer. Transformer++ (W=256) incorporates a window size of 256, retains the same model specification, and is also based on FlashAttention-2.
Based [2] is implemented according to the official open-source code333https://github.com/HazyResearch/based, with a hidden dimension of 768 and 15 layers for the small-scale version; among them, there are 3 layers of window attention with 12 attention heads and a window size of 128, and 3 layers of TaylorExp linear attention with 12 attention heads. Mamba-1 [16] is implemented based on the official open-source code444https://github.com/state-spaces/mamba, with a hidden dimension of 768, 24 layers and a state size of 16 for the small-scale version; the large-scale version has a hidden dimension of 2048, 48 layers and the same state size of 16. Mamba-2 [11] is implemented based on the official open-source code (URL is the same as Mamba-1), with a hidden dimension of 768, 24 layers and a state size of 128 for the small-scale version; the large-scale version has a hidden dimension of 2048, 48 layers and the same state size of 128. RWKV-5/6/7 [32, 33] are implemented based on the official open-source code555https://github.com/BlinkDL/RWKV-LM, with a hidden dimension of 768, 12 layers and 12 attention heads for the small-scale versions. DeltaNet [40] is based on Flash Linear Attention666https://github.com/fla-org/flash-linear-attention [52], with a hidden dimension of 768, 12 layers and 8 attention heads for the small-scale version. Gated DeltaNet [51] is also based on Flash Linear Attention, with a hidden dimension of 768, 12 layers, 6 attention heads and a head dimension of 96 for the small-scale version.
Top-1 Element is implemented using PyTorch777https://pytorch.org/ and the transformers888https://github.com/huggingface/transformers library, matching the model specifications of the Transformer++. Top-1 Chunk (Exact/Approx.) shares the same model specifications as Transformer++, with a chunk size of 128. Its sequence modeling part is implemented using Triton999https://github.com/triton-lang/triton [46], and its offset and index CUDA kernel is modified from the fastmoe101010https://github.com/laekov/fastmoe [18].
In addition, all of the above models use the Mixtral [22] vocabulary, which contains 32,000 tokens. Large-scale models involved in the training process were trained using the DeepSpeed111111https://github.com/deepspeedai/DeepSpeed open-source project. All models involving Rotary Embedding have a base of 10,000.
@@
Transformer++ is based on FlashAttention-2, while Transformer++ (W=256) uses FlexAttention121212https://pytorch.org/blog/flexattention/ [14]. The implementations of Top-1 Chunk (Exact) and Top-1 Chunk (Approx.) follow those described in the previous configuration. All four models have a hidden dimension of 768, 12 layers and 12 attention heads, with a chunk size of 128 for the two Top-1 Chunk models. Additionally, RWKV-7 was implemented using Flash Linear Attention when the sequence length was set to 100k to improve time efficiency. Gradient checkpointing was applied to reduce GPU memory consumption across all models. All models use the Mixtral vocabulary with a size of 32,000 tokens. Models incorporating Rotary Embedding maintain a base of 1,000,000.
@@ Appendix B Evaluation Procedures
For few-shot learning evaluation, general tasks include MMLU [19], OpenBookQA [29], ARC Easy & Challenge [9], BoolQ [8], RACE [24], SIQA [39], SCIQ [48], HellaSwag [55], COPA [37], PIQA [5], WinoGrande [38] and Winograd [25]; retrieval tasks include SWDE [26], SQuAD [36] and FDA [3]. Among these, MMLU uses a 5-shot setting, while all others are evaluated in a 0-shot setting. To ensure the stability and reproducibility of the evaluation results, we use the lm-evaluation-harness131313https://github.com/EleutherAI/lm-evaluation-harness [15] open-source evaluation framework for evaluation.
For time efficiency evaluation (Table 3), we compare the pre-training speed of different models with an input length of 100k tokens and a total batch size of 8. For inference speed comparison, we use an input length of 50k tokens and generate 5k tokens, with a total batch size of 64.
@@ Appendix C Related Work
Appendix D Algorithm Details
Specifically, we designed Triton kernels for Top-1 Chunk Selection architecture, with its forward pass detailed in Algorithm 1 and its backward pass outlined in Algorithm 2 . Beyond these two core algorithms, the implementation involves several other components. In both algorithms, and represent the offset and index, respectively. The offset sequentially records the number of times each chunk is selected by querys, while the index sequentially records the indices of all queries associated with each chunk. These can be obtained through a naive PyTorch implementation; we developed hardware-efficient CUDA kernels based on the logic of the naive implementation. Additionally, the chunk indices for both exact selection and approximate selection can also be acquired via a naive PyTorch implementation. Similarly, we developed hardware-efficient Triton kernels following the logic of the naive implementation. Additionally, we designed Triton kernels for the decoding stage, which can effectively accelerate the decoding process. The source code has been released at: https://github.com/luxinxyz/TSA .
Appendix E More Results Under Mixed Domain Pre-Training Setting
Although our primary focus is on experiments under the limited domain pre-training setting, we also conducted preliminary experiments on the Top-1 Element Selection Architecture under mixed domain pre-training setting during the early stages of this work. The results, as shown in Figures 8 and 9, indicate that the performance of the Top-1 Element Selection Architecture aligns with the conclusions drawn in the paper, with no degradation in base capabilities. This suggests that the usability of the proposed architecture remains consistent across different pre-training settings.